OpenStudy (anonymous):

Try this one: If \( X = 2^n \), how many \( n \), positive integer, exist such that \( X \) does not have a digit which is a power of \( 2 \)?

OpenStudy (anonymous):

The difficulty of the problem lies in the proving or disproving the number of numbers of such form existence or non-existence.

OpenStudy (anonymous):

\[1,2,4,8,16,32,64,128,256,512,1024,2048, \ldots\] Might help lol

OpenStudy (anonymous):

lol, Yes, but brute forcing won't give you any conclusive result though.

OpenStudy (anonymous):

Is the Binomial Theorem any help here?

OpenStudy (anonymous):

\[2^n = \sum_{k=0}^{n} \left(\begin{array}r n\\k \end{array}\right)\]

OpenStudy (mr.math):

\(2\) I think.